You borrow $1400. To repay the amount, you make 12 equal monthly payments of $122.50 Determine the following:
a) The effective monthly interest rate
b) The nominal annual interest rate
c) The effective annual interest rate
1
Expert's answer
2018-09-06T15:57:08-0400
You borrow $1400. To repay the amount, you make 12 equal monthly payments of $122.50.
Determine the following:
a) The effective monthly interest rate is:
r = (1 + i/n)^n - 1
i = 122.5*12/1400 - 1 = 0.05 or 5%.
r = (1 + 0.05/12)^12 - 1 = 0.0512 or 5.12%.
b) The nominal annual interest rate is:
i = 122.5*12/1400 - 1 = 0.05 or 5%.
c) The effective annual interest rate is:
r = (1 + 0.05/1)^1 - 1 = 0.05 or 5%.